OpenStudy (anonymous):

2'; 2. Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8)

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

first we need to find the slope

OpenStudy (anonymous):

so 8-5/6-4

OpenStudy (aripotta):

yes, which simplifies to...?

OpenStudy (anonymous):

3/2

OpenStudy (aripotta):

yes. now we can put it into point-slope form

OpenStudy (anonymous):

wait but all 4 answers are like this look

OpenStudy (aripotta):

yes, we have to put it into point-slope before we can put them into slope-intercept

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

so what's our point-slope form look like?

OpenStudy (aripotta):

you can use either given point

OpenStudy (anonymous):

m=3/2

OpenStudy (aripotta):

yea, that's the slope, but what does our equation look like

OpenStudy (anonymous):

im confuse about tht already lool

OpenStudy (aripotta):

we'll use the first point (4,5) so use that and m=3/2 and put it into point-slope form

OpenStudy (anonymous):

m=3/2+4

OpenStudy (aripotta):

no, into point-slope form...you forgot what that was? it's y - y1 = m(x - x1)

OpenStudy (anonymous):

y-5=m(x-4)

OpenStudy (aripotta):

yes, but we know that m = 3/2

OpenStudy (anonymous):

y=3/2x+1

OpenStudy (aripotta):

y = 3/2x - 1

OpenStudy (anonymous):

why -?

OpenStudy (aripotta):